A bird of mass 1 kg is flying due south at the latitude of 30 degree N in the northern hemisphere at a speed of 1 meter per second. Determine the coriolis force acting on it
1
Expert's answer
2012-10-08T09:54:26-0400
Fk=2*m*v*w*sin a
w=2Pi/24*3600 (rad/s)

v=1 m/s

m=1 kg

sin a=1 (orthogonality of vectors)

Fk=2*7.2*10^-5= 1.4*10^-4 N